OpenStudy (gigi1103):

Ken has $950 in a savings account at the beginning of the summer. He wants to have at least $600 in the account by the end of the summer. He withdraws $35 a week for food, clothes and movie tickets. Write an inequality for the number of weeks Ken can withdraw money, and describe the solution.

OpenStudy (gigi1103):

A) 950 - 35w ≥ 600; w ≤ 10 B) 950 - 35w ≥ 600; w ≤ 11 C) 950 + 35w ≥ 600; w ≥ 10 D) 950 + 35w ≥ 600; w ≥ 11

OpenStudy (gigi1103):

@Nnesha

OpenStudy (gigi1103):

@RainbowBrony555

OpenStudy (anonymous):

What do you think? :D

OpenStudy (gigi1103):

@dakid88

OpenStudy (dakid88):

B.

OpenStudy (anonymous):

-.-

OpenStudy (gigi1103):

thanks!!!

OpenStudy (gigi1103):

can you help with more

OpenStudy (dakid88):

yes

OpenStudy (gigi1103):

k

OpenStudy (gigi1103):

A bike rental company charges $50 for the first hour, and $25 for each additional hour. Michael must spend no more than $200 on his bike rental for the day. What is the maximum number of hours that Michael can rent a bike? A) 3 hours B) 4 hours C) 5 hours D) 7 hours

OpenStudy (gigi1103):

@dakid88

OpenStudy (dakid88):

Let x = number of hours (after that initial hour) and y = total cost so if "A bike rental company charges $50 for the first hour, and $25 for each additional hour.", then y = 25x + 50 ----------------------- We're told that "Michael must spend no more than $200", so y <= 200 y <= 200 25x + 50 <= 200 25x <= 200-50 25x <= 150 x <= 150/25 x <= 6 So the max number of hours after that first hour is 6 Add on the first hour to get 6+1 = 7 So the max number of hours total he can rent the bike is 7 hours.

OpenStudy (gigi1103):

thank you!! i have a few more is that ok?

OpenStudy (dakid88):

yep

OpenStudy (gigi1103):

k one sec

OpenStudy (gigi1103):

Which inequality's solution is graphed here? A) x = 0 B) x + 2 ≥ 0 C) x + 2 ≤ 0 D) x - 2 ≥ 0

OpenStudy (dakid88):

hold on

OpenStudy (dakid88):

im looking at the pic

OpenStudy (dakid88):

C.

OpenStudy (gigi1103):

k

OpenStudy (gigi1103):

Which inequality's solution is graphed here? A) x + 9 < 5 B) x - 5 < 9 C) x + 5 < 9 D) x - 9 < 5

OpenStudy (gigi1103):

i think c

OpenStudy (gigi1103):

wby?

OpenStudy (dakid88):

um i think c too

OpenStudy (gigi1103):

k

OpenStudy (gigi1103):

Which expression is equivalent to 11(17 – x)?

OpenStudy (gigi1103):

Which expression is equivalent to 11(17 – x)? A) 11 • 17 – 11x B) 11 + 17 – x C) 17 – 11x D) 11 • 17 – x

OpenStudy (gigi1103):

you threre

OpenStudy (gigi1103):

?